You are given the following information for Lightning Power Co. Assume the company’s tax rate is 23 percent. Debt: 13,000 6.2 percent coupon bonds outstanding, $1,000 par value, 28 years to maturity, selling for 108 percent of par; the bonds make semiannual payments. Common stock: 460,000 shares outstanding, selling for $64 per share; beta is 1.13. Preferred stock: 20,000 shares of 4 percent preferred stock outstanding, currently selling for $85 per share. The par value is $100 per share. Market: 6 percent market risk premium and 5.1 percent risk-free rate. What is the company's WACC?
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=64*460000
=29440000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*13000*1.08
=14040000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=85*20000
=1700000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=29440000+14040000+1700000
=45180000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 29440000/45180000
W(E)=0.6516
Weight of debt = MV of Bond/MV of firm
Weight of debt = 14040000/45180000
W(D)=0.3108
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 1700000/45180000
W(PE)=0.0376
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 5.1 + 1.13 * (6)
Cost of equity% = 11.88
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =28x2
1080 =∑ [(6.2*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^28x2
                   k=1
YTM = 5.6290145584
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.6290145584*(1-0.23)
= 4.334341209968
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 4/(85)*100
=4.71
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=4.33*0.3108+11.88*0.6516+4.71*0.0376
WACC =9.26%